Andrea's parents, a high school student, decided to give their daughter $ 100,000 for her postgraduate studies that she will study abroad in 5 years, and they also plan to save an additional $ 2,000 each month during this same time, since they estimate that their studies and support will cost approximately $ 1,500,000. The banking institution offers you to invest the $ 100,000 and the monthly savings in a financial instrument that provides you with a fixed nominal interest rate of 5% per year, compounded monthly.

What is the annual effective interest rate?

 [Value 2 points]

Select one:
a. 3.2525%
b. 5.1162%
c. 6.1212%
d. 4.1666%
Answer #1

Effective annual rate is calculated using the below formula:

EAR= (1+r/n)^n-1

Where r is the interest rate and n is the number of compounding periods in one year.

EAR= (1+0.05/12)^12 - 1

= 1.0512 - 1

= 0.0512*100

= 5.1162%.

Hence, the answer is option b.

In case of any query, kindly comment on the solution.